A firm wishes to issue new shares of its stock, which already trades in the market. The current stock price is $20, the most recent dividend was $2 per share, and the dividend is expected to grow at a rate of 4% forever. Flotation costs for this issue are expected to be 14%. What is the required rate of return (or financing cost) in this new issue?

Note: when flotation costs are given as a percentage instead of in dollar terms, the denominator in the formula changes from (P-F) to P*(1-F).

Enter your answer as a percentage, rounded to two decimals. So, if your answer is 0.123456, enter 12.34.
Answer #1

required rate of return = (next year dividend / current share price net of flotation costs) + constant growth rate

next year dividend = $2 + 4% = $2.08

current share price net of flotation costs = $20 - 14% = $17.20

required rate of return = (2.08 / 17.20) + 4% = 16.09%